Write a quadratic equation and solve
Find two integers such that the first is 2 more than twice the second and their product is 112.

Answer by mananth(16946)


Find two integers such that the first is 2 more than twice the second and their product is 112.
let second number be x
first number => 2x+2
x(2x+2) = 112
3x^2+2x=112
/2
x^2+x=56
x^2+x-56=0
x^2+8x-7x-56=0
x(x+8)-7(x+8)=0
(x+8)(x-7)=0
x=-8 OR x=7
2x+2 when x=-8
=-14
(-8,-14)
2x+2 when x=7
=16
(7,16)